OpenStudy (anonymous):

solve equation by elimination 8u_3v=-16 3u+v=-8

OpenStudy (anonymous):

supposed to be 8u-3v=-16

OpenStudy (anonymous):

u=-40/17, v=-16/17 via substitution method